# Can Large Language Models Detect Misinformation in Scientific News Reporting?

### v1

@@ Abstract.
Scientific facts are often spun in the popular press with the intent to influence public opinion and action, as was evidenced during the COVID-19 pandemic. Automatic detection of misinformation in the scientific domain is challenging because of the distinct styles of writing in these two media types and is still in its nascence. Most research on the validity of scientific reporting treats this problem as a claim verification challenge. In doing so, significant expert human effort is required to generate appropriate claims. Our solution bypasses this step and addresses a more real-world scenario where such explicit, labeled claims may not be available. The central research question of this paper is whether it is possible to use large language models (LLMs) to detect misinformation in scientific reporting. To this end, we first present a new labeled dataset SciNews, containing 2.4k scientific news stories drawn from trusted and untrustworthy sources, paired with related abstracts from the CORD-19 database, a comprehensive collection of COVID-19 and historical coronavirus research. Our dataset includes both human-written and LLM-generated news articles, making it more comprehensive in terms of capturing the growing trend of using LLMs to generate popular press articles. Then, we identify dimensions of scientific validity in science news articles and explore how this can be integrated into the automated detection of scientific misinformation. We propose several baseline architectures using LLMs to automatically detect false representations of scientific findings in the popular press. Our first architecture, SERIf, uses three modules: Summarization, Evidence Retrieval, and Inference to classify the news article as fake or true; the second architecture, SIf, bypasses the explicit evidence retrieval module while keeping the other two, and the third, direct-to-inference architecture, D2I, where both the summarization and the explicit evidence retrieval module are removed. For each of these architectures, we use several prompt engineering strategies including: zero-shot, few-shot, and chain-of-thought prompting. We also test these architectures and prompting strategies on GPT-3.5, GPT-4, and Llama2-7B, Llama2-13B. We note that the SIf architecture does best while the naive D2I does worst in terms of accuracy and F-1 scores.
@@ 1. Introduction
Scientific information is communicated to the non-expert audience via popular press (news articles) and online platforms like blogs, social media posts, etc. Studies have shown that news with scientific-sounding content is trusted more than other types (Sharon and Baram-Tsabari, 2020). Therefore, any misinformation in the scientific domain can cause significant public risk as was evidenced during the recent COVID-19 pandemic (Mheidly and Fares, 2020). Other examples include the emergence of vaccination hesitancy (Razai et al., 2021; Baines et al., 2021), eroded trust in health institutions (Rodriguez-Morales and Franco, 2021), and the amplification of public fear and anxiety (Nelson et al., 2020; Verma et al., 2022). Therefore, it is imperative to identify misinformation in scientific news reporting.
Several websites are maintained by science reporters (Health News Review [^1] https://www.healthnewsreview.org/) and scientists (Science Feedback  [^2] https://science.feedback.org/) to track scientific misinformation in the media. Although such manual debunking is important, the sheer volume of scientific news can make this task unscalable.
Natural language processing (NLP) based approaches have consequently started to emerge to deal with this problem. These methods typically involve language analysis, like detecting exaggeration (Wright and Augenstein, 2021a) and certainty (Pei and Jurgens, 2021) and fact-checking (Guo et al., 2022a) and claim verification (Pradeep et al., 2020). Several claim verification datasets have also been developed for this problem (Thorne et al., 2018; Wadden et al., 2022) and a method for modeling information change from scientific article to scientific reporting has also been proposed (Wright et al., 2022a).
@@
To answer the above questions, we first create a novel dataset, SciNews [^3] We will make our dataset fully accessible to the research community, comprised of scientific news and related scientific articles. Given the rising trend in LLM-generated content in both legitimate reporting and misinformation, this dataset contains an equal number of LLM-generated and human-written articles. Our dataset construction method does away with extensive human annotation which can prove to be very expensive. The dataset construction pipeline is flexible enough to allow continuous updates with emerging news articles and scientific articles.
We then propose three architectures using LLMs to automatically detect false representations of scientific findings in the popular press without explicit claim generation. The first architecture, SERIf, uses three modules: Summarization, Evidence Retrieval, and Inference to classify the news article as fake or true; the second architecture, SIf, bypasses the explicit evidence retrieval module while keeping the other two, and the third, direct-to-inference architecture, D2I, adopts a direct-to-inference approach, dispensing with both summarization and explicit evidence retrieval. For each of these architectures, we employ several prompt engineering strategies including zero-shot, few-shot, and chain-of-thought prompting. We test these architectures using several state-of-the-art LLMs, including GPT (versions GPT-3.5 and GPT-4) and Llama2 (models with 7B and 13B).
Our extensive experiments conducted on the SciNews dataset have yielded significant findings: 1) we discovered that false scientific content generated by large language models (LLMs) is more challenging to detect compared to misinformation written by humans. 2) the SIf architecture arises capable of matching, and in some cases, surpassing the performance of the SERIf architecture. This suggests the feasibility of using LLMs to develop a more generalized architecture for detecting scientific misinformation more appropriate for real-life situations. 3) We make the first attempt to define dimensions of scientific validity and incorporate them into the Chain-of-Thought prompt. This novel addition has been instrumental in enhancing the LLMs’ ability to make more accurate predictions. 4) LLMs can generate plausible explanations for their decision-making process, which can help users understand the model’s judgment of news veracity.
@@ 2.1. Fact-Checking
Automatic fact-checking, which assesses the truthfulness of claims made in the text (Vlachos and Riedel, 2014; Guo et al., 2022b), has been extensively studied across various domains, including common knowledge verification (Thorne et al., 2018), political topics (Wang, 2017), COVID-19 (Saakyan et al., 2021), E-commercial (Zhang et al., 2020), biology (Chen et al., 2022). When thinking about scientific misrepresentation in popular media, it is natural to think of the veracity of scientific findings. It is therefore possible to cast the scientific misinformation as a fact-checking problem or claim verification problem. Several researchers have taken this approach to defect misinformation in the scientific domain (Vladika and Matthes, 2023; Wadden et al., 2020; Diggelmann et al., 2020; Wadden and Lo, 2021; Wadden et al., 2022; Wright et al., 2022b; Wang et al., 2023). These works typically construct claims from the existing scientific literature by manually reformulated scientific findings and then the constructed claims are verified by utilizing pre-existing knowledge resources. However, most of these works rely on human resources to identify and extract appropriate claims for verification. Furthermore, these artificially constructed claims may not accurately represent the complexity and nuance of claims encountered in real-world scenarios.
@@ 2.2. Scientific Language Analysis
Scholarly document processing has garnered considerable attention in recent years, reflecting a growing interest in the analysis and interpretation of scientific literature (Chandrasekaran et al., 2020). Of particular relevance to our research are tasks that track the change of scientific information from published literature to social press. This includes investigating writing strategies employed in science communication  (Tan and Lee, 2014; August et al., 2020), detecting changes in certainty (Pei and Jurgens, 2021) and exaggeration detection (Wright and Augenstein, 2021b). Furthermore, the automatic detection of semantic similarities between scientific texts and their paraphrases represents an alternative approach for analyzing scientific content (Lo et al., 2019; Piskorski et al., 2023). However, rather than use the typical metrics for measuring semantic similarity, we propose to use the inherent knowledge in pre-trained LLMs for this task.
@@ 2.3. Large Language Model in Misinformation
LLMs have consistently demonstrated the ability to generate text on par with human authors  (Zhang et al., 2023a; Yang et al., 2023). This has led to their widespread use by professionals in generating legitimate real news stories. Unfortunately, they have also been used to generate misinformation (Zhou et al., 2023; Chen and Shu, 2023; Buchanan et al., 2021) and often at a much larger scale than is humanly possible. While falsehoods crafted by LLMs prove challenging for humans to detect, compared to human-generated ones (Chen and Shu, 2023), several studies have illustrated the feasibility of identifying LLM-generated text using  (Tang et al., 2023; Buchanan et al., 2021). Motivated by these studies, we include a balanced set of LLM-generated scientific articles, both fake and true, in the SciNews dataset.
3. SciNews Dataset Construction
@@
The first part of this work is to create a dataset, SciNews, of scientific news articles and associated technical scientific publications. The SciNews Dataset contains 2,400 news articles which are labeled: Reliable or Unreliable depending on whether the article represents scientific fact truthfully or not, respectively. The dataset contains 1,200 articles in each category to keep it balanced. Each article is systematically paired with up to three pertinent scientific abstracts from the CORD-19 repository. The CORD-19 is a comprehensive resource of over 1 million scholarly articles, including over 300,000 with full text, about COVID-19, SARS-CoV-2, and related coronaviruses (Wang et al., 2020).
Given the growing trend in using Large Language Models (LLMs) to generate legitimate as well as fake news stories (Zhou et al., 2023; Chen and Shu, 2023), we include LLM-generated news articles in the dataset. The SciNews dataset contains an equal number (1,200) of human-written and LLM-generated news articles in each category (reliable and unreliable). The statistics of the SciNews dataset are presented in Table 1. We show the construction of SciNews in Figure 1 and describe the construction of the dataset below.
Table 1. Distribution of article labels in SciNews Dataset.
@@ 3.1.1. Leveraging Publicly Available Dataset
We leveraged MMCoVaR (Chen et al., 2021), COVID19-FNIR (Saenz et al., 2021), and COVID-Rumor (Cheng et al., 2021), which are labeled datasets containing human written news articles on COVID-19 from January 2020 through May 2021.
@@ 3.2.2. Jailbreak Prompts
We utilized the curated collection of abstracts as guiding resources for LLMs to generate both reliable and unreliable scientific articles. To mitigate the risk of generating harmful content, LLMs are subjected to an alignment process, complemented by the setup of predefined prompts serving as security measures (Achiam et al., 2023; Qi et al., 2023; TermsOfUseBing, 2023). Therefore, requests for LLMs to generate fake messages are usually denied. In light of these constraints and inspired by the Jailbreak Attacks (Mowshowitz, [n. d.]; Liu et al., 2023b, a; Shen et al., 2023), we designed a ‘Jailbreak’ prompt to enable the LLMs to generate fake news-oriented scientific articles that were both informative and contextually aligned with the provided abstracts. The designed jailbreak prompt is illustrated in Figure 2.
@@
Given selected scientific abstract resources, we simulated a scenario where we acted as the instructor of a science class, prompting the large language model (LLM) to create two types of articles: a ‘True Article’ and a ‘Convincing False Article’. The objective was to use these articles as teaching tools to help students discern between authentic scientific knowledge and fabricated information. By including both true and false LLM-created content we can ensure that the systems trained on our dataset to detect false content are not simply detecting LLM-generated content, but would be adept at distinguishing between true and false content. The generated articles are generally in the style of a news article, with many including an explicit title, to mimic human-generated scientific news. In consideration of the cost implications associated with different LLMs [^5] https://openai.com/pricing666llama.meta.com, we primarily utilized Llama2-7B for generating the bulk of data samples, supplemented by a smaller set from GPT-3.5. After filtering (see subsection 3.4), this approach led to the creation of 1,000 data samples from Llama2-7b and 200 from GPT-3.5, reflecting a balance between quality and resource optimization.
@@ 3.4. Quality Control
For the collection of human-written news articles from various sources, we referred to the guidelines outlined in  (Pustejovsky and Stubbs, 2012). Based on its principles and our specific needs, we developed an instruction guide, which can be found in Appendix A.5. To ensure uniformity and understanding of the task, all team members thoroughly reviewed this guide. An additional layer of quality assurance involved cross-checking the collected data among team members. This step was implemented to mitigate any potential biases and to guarantee that the data aligned with our collection criteria.
@@ 4.2.2. Summarization module
Inspired by the recent success of text summarization (Zhang et al., 2023b) using LLMs, we take the ’Extractive - Abstractive’ two-step summarization strategy (Zhang et al., 2023c) to construct a summary of the news article. The extractive summarization process summarizes the article by identifying and concatenating the most salient sentences from it, ensuring that the extracted summaries are consistent with the original text. The resulting summaries serve as a foundation for the abstractive summarization which uses a generative approach to create a more concise and cohesive summary. By synthesizing the extractive and abstractive approaches, the module ensures a balance between accuracy (adherence to the original text) and brevity (conciseness and essence of the content), thereby providing an effective and reliable summarization for further analysis in the misinformation detection process.
@@
As a sanity check, to verify the quality of the summary, we randomly selected 200 samples for annotation by two graduate students with a background in NLP, and evaluated the summaries using four criteria: 1) quality of extractive summary (High vs Low); 2) quality of abstractive summary (High vs Low); 3) presence of hallucination in abstractive summary (Yes vs No); 4) comparison of Extractive and Abstractive Summaries. We used the Krippendorff’s alpha score to evaluate the agreement between the annotations (Hayes and Krippendorff, 2007). The alpha scores for the four aspects were 0.53, 0.82, 0.91, and 0.94, respectively. These values suggest that there is strong agreement that the abstractive summaries effectively encapsulate the core information of the original texts and maintain a high degree of consistency. Consequently, we have decided to use the abstractive summaries in subsequent steps of our analysis.
@@ 4.2.4. Inference Module
The Inference Module is dedicated to assessing the the veracity of the summarized news paragraph using the set of retrieved evidence sentences using the abstractive summary, $S_{a}$, and the selected evidence sentence set $\{e_{i}\}$. Thus, the inference module produces a binary output (reliable or unreliable) for each $<S_{a},\{e_{i}\}>$ pair.
Table 2. Performance results of our proposed SERIf architecture using different LLMs and prompt strategies.
Models Prompt Strategy Human-Written LLM-Generated Overall
@@
GPT-3.5 Zero-Shot 74.25 74.95 72.83 73.87 66.75 60.88 93.66 73.80 70.50 67.92 83.23 73.84
@@
CoT 76.67 76.20 77.67 76.92 66.75 60.66 95.33 74.14 71.71 68.43 86.50 75.53
GPT-4 Zero-Shot 77.33 76.48 79.00 77.72 70.25 62.91 98.67 76.83 73.79 69.70 88.34 77.26
@@
CoT 79.58 76.90 85.00 80.72 67.25 60.50 99.33 75.20 73.42 68.70 92.17 77.96
LLAMA2-7B Zero-Shot 56.00 53.24 98.33 69.10 51.17 50.60 96.80 66.50 53.89 51.92 97.57 67.80
@@
CoT 56.83 59.20 97.80 73.70 51.58 50.80 96.80 66.70 54.21 55.00 97.30 70.20
LLAMA2-13B Zero-Shot 57.33 59.50 98.50 74.20 53.67 57.10 97.00 71.90 55.50 58.30 97.75 73.05
@@
CoT 58.00 59.90 99.00 74.60 55.08 52.70 98.30 68.60 56.54 56.30 98.65 71.60

### v2

@@ Abstract.
Automatic detection of misinformation in the scientific domain is challenging because of the distinct styles of writing in scientific publications vs reporting. This problem is exacerbated by the prevalence of large language model generated misinformation. In this paper, we address the problem of automatic detection of misinformation in a more realistic scenario where there is no prior knowledge of the origin (LLM or human written) of the text, and explicit claims may not be available. We first introduce a novel labeled dataset, CoSMis (SciNews), comprising of 2,400 scientific news stories sourced from both reliable and unreliable outlets, paired with relevant abstracts from the CORD-19 database. Our dataset uniquely includes both human-written and LLM-generated news articles. We propose a set of dimensions of scientific validity (DoV) along which to evaluate the articles for misinformation. These are then incorporated into the prompt structures for the LLMs. We propose three LLM pipelines to compare scientific news to relevant research papers and classify for misinformation. The three pipelines represent different levels of intermediate processing steps on the raw scientific news articles and research papers. We apply various prompt engineering strategies: zero-shot, few-shot, and DoV-guided Chain-of-Thought prompting, to these architectures and evaluate them using GPT-3.5, GPT-4, Llama2-7B/13B/70B and Llama3-8B. [^1] This work has been accepted by AAAI-25 Workshop on Preventing and Detecting LLM Misinformation (Link). [^2] We release the dataset at: https://github.com/InfintyLab/CoSMis-SciNews-
@@ 1. Introduction
Scientific information is communicated to the non-expert audience via popular press (news articles) and online platforms like blogs, social media posts, etc. Studies have shown that news with scientific-sounding content is trusted more than other types (ShaEtal20). Therefore, any misinformation in the scientific domain can cause significant public risk as was evidenced during the recent COVID-19 pandemic (MheEtal20). Other examples include the emergence of vaccination hesitancy (RazEtal21; BaiEtal21), eroded trust in health institutions (RodEtal21), and the amplification of public fear and anxiety (NelEtal20; VerEtal22). Therefore, it is imperative to identify misinformation in scientific news reporting.
Several websites are maintained by science reporters (Health News Review [^3] https://www.healthnewsreview.org/) and scientists (Science Feedback  [^4] https://science.feedback.org/) to track scientific misinformation in the media. Although such manual debunking is important, the sheer volume of scientific news can make this task unscalable. Natural language processing (NLP) based approaches have consequently started to emerge to deal with this problem. These methods typically involve language analysis, like detecting exaggeration (WriEtal21) and certainty (PeiEtal21) and fact-checking (ZhiEtal22) and claim verification (RonEtal20). Several claim verification datasets have also been developed for this problem (ThoEtal18; WaEtal22) and a method for modeling information change from scientific article to scientific reporting has also been proposed (DusEtal22).
@@
To answer the above questions, we first create a novel COVID related Scientific Misinformation (CoSMis, or SciNews [^5] https://github.com/InfintyLab/CoSMis-SciNews-) dataset, comprised of scientific news and related scientific articles. Given the rising trend in LLM-generated content in both legitimate reporting and misinformation, this dataset contains an equal number of LLM-generated and human-written articles. The dataset construction pipeline is flexible enough to allow continuous updates with emerging news articles and scientific articles.
We then propose three architectures using LLMs to automatically detect false representations of scientific findings in the popular press without explicit claim generation. The first architecture, SERIf, uses three modules: Summarization, Evidence Retrieval, and Inference to classify the news article as fake or true; the second architecture, SIf, bypasses the explicit evidence retrieval module while keeping the other two, and the third, direct-to-inference architecture, D2I, adopts a direct-to-inference approach, dispensing with both summarization and explicit evidence retrieval. For each of these architectures, we employ several prompt engineering strategies including zero-shot, few-shot, and chain-of-thought prompting. We test these architectures using several state-of-the-art LLMs, including GPT (3.5&4), Llama2 (7B,13B&70B) and Llama3 (8B).
This work makes the following contributions: 1) introducing CoSMis (SciNews), a unique dataset designed for detecting scientific misinformation, which includes human-authored articles and LLM-generated texts to mirror real-world challenges. 2) proposing three LLM pipelines to detect scientific misinformation “in the wild” using scientific articles as grounding evidence material. 3) proposing Dimensions of Validity (DoV) guided chain-of-thought prompting 4) testing the proposed pipelines on the architectures on the CoSMis dataset and demonstrating that LLMs are able to detect scientific misinformation without needing a training phase and testing phase and 5) demonstrating that the DoV prompting can be used to derive explanations for the LLM’s decision.
@@ 2.1. Fact-Checking
Automatic fact-checking, which assesses the truthfulness of claims made in the text (VlaRie14; GuoEtal22), has been extensively studied across various domains, including common knowledge verification (ThoEtal18), political topics (Wan17), COVID-19 (SaChMu21), E-commercial (ZhaEtal20E), biology (WenEtal22b). When thinking about scientific misrepresentation in popular media, it is natural to think of the veracity of scientific findings. It is therefore possible to cast the scientific misinformation as a fact-checking problem or claim verification problem. Several researchers have taken this approach to defect misinformation in the scientific domain (VlaMat23; WaEtal20; ThoEtal20; WadKyl21; WaEtal22; WriEtal22; WaEtal23). These works typically construct claims from the existing scientific literature by manually reformulated scientific findings and then the constructed claims are verified by utilizing pre-existing knowledge resources. However, most of these works rely on human resources to identify and extract appropriate claims for verification. Furthermore, these artificially constructed claims may not accurately represent the complexity and nuance of claims encountered in real-world scenarios.
@@ 2.2. Scientific Language Analysis
Scholarly document processing has garnered considerable attention in recent years, reflecting a growing interest in the analysis and interpretation of scientific literature (ChaEtal20overview). Of particular relevance to our research are tasks that track the change of scientific information from published literature to social press. This includes investigating writing strategies employed in science communication  (TanLee14; AugEtal20), detecting changes in certainty (PeiEtal21) and exaggeration detection (WriIsa21). Furthermore, the automatic detection of semantic similarities between scientific texts and their paraphrases represents an alternative approach for analyzing scientific content (KylEtak19; JakEtal23). However, rather than use the typical metrics for measuring semantic similarity, we propose to use the inherent knowledge in pre-trained LLMs for this task.
@@ 2.3. Large Language Model in Misinformation
LLMs have consistently demonstrated the ability to generate text on par with human authors  (ZhaEtal23; YanEtal23). This has led to their widespread use by professionals in generating legitimate real news stories. Unfortunately, they have also been used to generate misinformation (ZhoEtal23; CheShu23; Buc21Etal) and often at a much larger scale than is humanly possible. While falsehoods crafted by LLMs prove challenging for humans to detect, compared to human-generated ones (CheShu23), several studies have illustrated the feasibility of identifying LLM-generated text using  (TanChuHu23; Buc21Etal). Motivated by these studies, we include a balanced set of LLM-generated scientific articles, both fake and true, in the SciNews (CoSMis)dataset.
3. CosMIs Dataset Construction
@@
The first part of this work is to create a dataset, CoSMIs (SciNews), of scientific news articles and associated technical scientific publications. The CoSMIs Dataset contains 2,400 news articles which are labeled: Reliable or Unreliable depending on whether the article represents scientific fact truthfully or not, respectively. The dataset contains 1,200 articles in each category to keep it balanced. Each article is systematically paired with up to three pertinent scientific abstracts from the CORD-19 repository. The CORD-19 is a comprehensive resource of over 1 million scholarly articles, including over 300,000 with full text, about COVID-19, SARS-CoV-2, and related coronaviruses (Wan2Etal20). Given the growing trend in using Large Language Models (LLMs) to generate legitimate as well as fake news stories (ZhoEtal23; CheShu23), we include LLM-generated news articles in the dataset. The CoSMIs dataset contains an equal number (1,200) of human-written and LLM-generated news articles in each category (reliable and unreliable). The statistics of the CoSMIs dataset are presented in Table 1. We show the construction of CoSMIs in Figure 1 and describe the construction of the dataset below. We present more statistics in Appendix B.
Table 1. Distribution of article labels in CoSMIs (SciNews) Dataset.
@@ 3.1.1. Leveraging Publicly Available Dataset
We leveraged MMCoVaR (CheChuSub21), COVID19-FNIR (JulEtal21), and COVID-Rumor (MinEtal21), which are labeled datasets containing human written news articles on COVID-19 from January 2020 through May 2021.
@@ 3.2.2. Jailbreak Prompts
We utilized the curated collection of abstracts as guiding resources for LLMs to generate both reliable and unreliable scientific articles. To mitigate the risk of generating harmful content, LLMs are subjected to an alignment process, complemented by the setup of predefined prompts serving as security measures (ope23; QiEtal23; Bin23). Therefore, requests for LLMs to generate fake messages are usually denied. In light of these constraints and inspired by the Jailbreak Attacks (Mow22; LiuEtal23a; LiuEtal23b; SheEtal23), we designed a ‘Jailbreak’ prompt to enable the LLMs to generate fake news-oriented scientific articles that were both informative and contextually aligned with the provided abstracts. The designed jailbreak prompt is illustrated in Figure 2.
@@
Given selected scientific abstract resources, we simulated a scenario where we acted as the instructor of a science class, prompting the large language model (LLM) to create two types of articles: a ‘True Article’ and a ‘Convincing False Article’. The objective was to use these articles as teaching tools to help students discern between authentic scientific knowledge and fabricated information. By including both true and false LLM-created content we can ensure that the systems trained on our dataset to detect false content are not simply detecting LLM-generated content, but would be adept at distinguishing between true and false content. The generated articles are generally in the style of a news article, with many including an explicit title, to mimic human-generated scientific news. In consideration of the cost implications associated with different LLMs [^7] https://openai.com/pricing888llama.meta.com, we primarily utilized Llama2-7B for generating the bulk of data samples, supplemented by a smaller set from GPT-3.5. After filtering (see subsection 3.4), this approach led to the creation of 1,000 data samples from Llama2-7b and 200 from GPT-3.5, reflecting a balance between quality and resource optimization. We show the LLM-generated article example in Appendix C.2.
@@ 3.4. Quality Control
For the collection of human-written news articles from various sources, we referred to the guidelines outlined in  (JamAmb12). Based on its principles and our specific needs, we developed an instruction guide, which can be found in Appendix A.4. To ensure uniformity and understanding of the task, all team members thoroughly reviewed this guide. An additional layer of quality assurance involved cross-checking the collected data among team members. This step was implemented to mitigate any potential biases and to guarantee that the data aligned with our collection criteria.
@@ 4.2.2. Summarization module
Inspired by the recent success of text summarization (ZhaLadDurEtal23) using LLMs, we take the ’Extractive - Abstractive’ two-step summarization strategy (ZhaLiuZha23) to construct a summary of the news article. The extractive summarization process summarizes the article by identifying and concatenating the most salient sentences from it, ensuring that the extracted summaries are consistent with the original text. The resulting summaries serve as a foundation for the abstractive summarization which uses a generative approach to create a more concise and cohesive summary. By synthesizing the extractive and abstractive approaches, the module ensures a balance between accuracy (adherence to the original text) and brevity (conciseness and essence of the content), thereby providing an effective and reliable summarization for further analysis in the misinformation detection process.
@@
As a sanity check, to verify the quality of the summary, we randomly selected 200 samples for annotation by two graduate students with a background in NLP, and evaluated the summaries using four criteria: 1) quality of extractive summary (High vs Low); 2) quality of abstractive summary (High vs Low); 3) presence of hallucination in abstractive summary (Yes vs No); 4) comparison of Extractive and Abstractive Summaries. We used the Krippendorff’s alpha score to evaluate the agreement between the annotations (AndKla07). The alpha scores for the four aspects were 0.53, 0.82, 0.91, and 0.94, respectively. These values suggest that there is strong agreement that the abstractive summaries effectively encapsulate the core information of the original texts and maintain a high degree of consistency. Consequently, we have decided to use the abstractive summaries in subsequent steps of our analysis.
@@ 4.2.4. Inference Module
The Inference Module is dedicated to assessing the the veracity of the summarized news paragraph using the set of retrieved evidence sentences using the abstractive summary, $S_{a}$, and the selected evidence sentence set $\{e_{i}\}$. Thus, the inference module produces a binary output (reliable or unreliable) for each $S_{a},\{e_{i}\}>$ pair.
Table 2. Performance results of our proposed three architectures using different LLMs and different prompt strategies.
Models Arch Prompt Strategy Human-Written LLM-Generated Overall
@@
Proprietary Models
GPT3.5 SERIf Zero-Shot 74.25 74.95 72.83 73.87 66.75 60.88 93.66 73.80 70.50 67.92 83.23 73.84
@@
DoV-CoT 76.67 76.20 77.67 76.92 66.75 60.66 95.33 74.14 71.71 68.43 86.50 75.53
SIf Zero-Shot 78.67 79.76 76.83 78.27 62.00 57.00 99.00 72.00 70.34 68.38 87.92 75.14
Few-Shot 76.08 79.88 73.67 75.68 65.33 59.89 92.83 72.81 70.71 69.89 83.25 74.25
DoV-CoT 79.92 80.88 78.33 79.50 70.17 65.12 86.83 74.43 75.05 73.00 82.58 76.97
DoV-CoT 77.17 69.50 96.83 80.91 64.08 57.63 99.65 73.03 70.63 63.57 98.24 76.97
GPT-4 SERIf Zero-Shot 77.33 76.48 79.00 77.72 70.25 62.91 98.67 76.83 73.79 69.70 88.34 77.26
@@
DoV-CoT 79.58 76.90 85.00 80.72 67.25 60.50 99.33 75.20 73.42 68.70 92.17 77.96
SIf Zero-Shot 78.33 84.00 70.00 76.36 71.08 65.79 87.83 75.23 74.71 74.90 78.92 75.80
DoV-CoT 80.00 80.00 79.00 80.00 71.00 64.00 98.00 77.00 75.50 72.00 88.50 78.50
DoV-CoT 78.08 84.60 68.67 75.80 72.00 65.20 98.50 78.30 75.04 74.90 83.56 77.05
Open-Source Models
LLAMA2-7B SERIf Zero-Shot 56.00 53.24 98.33 69.10 51.17 50.60 96.80 66.50 53.89 51.92 97.57 67.80
@@
DoV-CoT 56.83 59.20 97.80 73.70 51.58 50.80 96.80 66.70 54.21 55.00 97.30 70.20
LLAMA2-13B SERIf Zero-Shot 57.33 59.50 98.50 74.20 53.58 52.80 97.20 68.40 55.46 56.15 97.85 71.30
@@
DoV-CoT 58.00 59.90 99.00 74.60 55.08 52.70 98.30 68.60 56.54 56.30 98.65 71.60
LLAMA2-70B SERIf Zero-Shot 67.08 60.04 99.00 75.00 55.00 52.73 96.67 68.30 61.04 56.39 97.84 71.65
DoV-CoT 67.50 60.82 98.33 75.12 53.67 57.10 97.00 71.90 60.59 58.96 97.67 73.51
LLAMA3-8B SERIf Zero-Shot 62.83 57.49 98.33 72.50 52.42 51.26 97.17 67.82 57.63 54.38 97.75 70.16
Few-Shot 53.67 51.98 97.52 71.21 52.42 51.26 97.17 67.13 53.05 51.62 97.35 69.17
DoV-CoT 65.25 59.31 97.17 73.68 56.25 53.42 97.67 68.96 60.75 56.37 97.42 71.32